OpenStudy (anonymous):

what is the mass in grams of 2.90x10^12 lead (Pb) in atoms ?

OpenStudy (mertsj):

What is the atomic weight of Pb?

OpenStudy (anonymous):

207.2

OpenStudy (mertsj):

\[2.90\times 10^{12}atoms Pb \times\frac{1 mole Pb}{6.023\times 10^{23}atoms Pb}\times\frac{207.2 g Pb}{1 mole Pb}\]

OpenStudy (mertsj):

I think that works out to 9.98 x10^-10

OpenStudy (anonymous):

how do you figure out the correct scientific notation ?

OpenStudy (mertsj):

I just plug it into my calculator

OpenStudy (mertsj):

But if you can't do that then do this:

OpenStudy (mertsj):

\[\frac{2.90(207.2)}{6.023}=99.8\]

OpenStudy (mertsj):

Now we must do the powers of 10

OpenStudy (mertsj):

\[\frac{10^{12}}{10^{23}}=10^{-11}\]

OpenStudy (mertsj):

So now we have: \[99.8\times 10^{-11}\]

OpenStudy (anonymous):

thank you !!

OpenStudy (mertsj):

And we need to put that into scientific notation. Can you do that?

OpenStudy (anonymous):

yes sir/ma'am

OpenStudy (mertsj):

Good for you.